OpenStudy (anonymous):

The path of a ball can be modelled with the function f(t) = -½g t² + vot + h, where g represents the force of gravity, t represents time, vo represents the initial velocity, and h represent the height from which the ball is thrown. Willie throws a basketball from the height of 1.8 m at a speed of 12 m/s. Write the equation that models this throw, given that Earth's gravity is 9.8. What will be the maximum height of the basketball rounded to 2 decimal places? How long before the basketball will be at 3 m to drop into the basket?

OpenStudy (phi):

Willie throws a basketball from the height of 1.8 m at a speed of 12 m/s. **Write the equation that models this throw, given that Earth's gravity is 9.8. that means replace g with 9.8 and replace vo with 12 , and replace h with 1.8 can you do that ?

OpenStudy (anonymous):

yea, how would i figure out the maximum height, and how long before the basketball will be at 3m?

OpenStudy (phi):

what did you get for the equation ?

OpenStudy (anonymous):

f(t) = -½(9.8) t² + 12 + 1.8 would i do -b/2a to find the vertex (maximum height)? Thanks for the quick reply

OpenStudy (phi):

I assume you mean f(t) = -½(9.8) t² + 12t + 1.8 (with 12t not just 12) yes you can find the t where it peaks. see http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php but you already know the formula a= -9.8/2 and 2a is -9.8 t= -b/2a = -12/-9.8 = 1.2245 seconds now use that value of t in the equation to find the height.

OpenStudy (phi):

**How long before the basketball will be at 3 m to drop into the basket? I can't answer this one. where is the basket ?

OpenStudy (anonymous):

It doesn't say :/ the only information given is stated in the question

OpenStudy (anonymous):

does it mean to substitute f(t) as 3?

OpenStudy (phi):

That is a good guess. I would answer that, and ask the teacher why they ask such vague questions.

OpenStudy (anonymous):

Haha, thanks very much for your help!